OpenStudy (anonymous):

If y =12 when x= 6, find x when y=-24

OpenStudy (anonymous):

find a point in a graph?

OpenStudy (freckles):

You would tell whoever gave this question that there isn't enough info.

OpenStudy (freckles):

does y vary directly with x maybe?

OpenStudy (freckles):

does y vary indirectly with x? or you not given anything else?

OpenStudy (anonymous):

yes, i want to know too..

OpenStudy (anonymous):

Write an inverse variation equation that relates x and y. Assume that y varies inversely as x. Then solve.

OpenStudy (freckles):

\[y=\frac{k}{x}\]

OpenStudy (freckles):

plug in x=6,y=12 to find the constant k

OpenStudy (freckles):

that is the first step then there is another step afterwards

OpenStudy (anonymous):

simplify and substitute

OpenStudy (anonymous):

okay i plugged everything in..

OpenStudy (freckles):

did you find the constant k?

OpenStudy (freckles):

that was the whole point in pluggin in your first part (x,y)=(6,12)

OpenStudy (freckles):

first pair*

OpenStudy (anonymous):

so it'd be -3 ?

OpenStudy (anonymous):

the answer is -12?

OpenStudy (anonymous):

y=kx 12=k(6) 2=k y=2x -24=2x x=-24/2 x=-12

OpenStudy (freckles):

I think 6(12) isn't -3

OpenStudy (anonymous):

thats not even a choice.

OpenStudy (freckles):

\[y=\frac{k}{x} \\ xy=k \\ 6(12) =k\]

OpenStudy (freckles):

do you know how to multiply 6 and 12?

OpenStudy (freckles):

if you don't you can think of it as 12+12+12+12+12+12 (the sum of six 12's)

OpenStudy (anonymous):

yes i do thanks but my teacher checked it and she said its right

OpenStudy (freckles):

oh you already did the question?

OpenStudy (anonymous):

i did what you said to start off with but then she emailed me back and set it up

OpenStudy (freckles):

ok so when you said -3 you weren't talking about k you were talking about the for x when y=-24

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[xy=k \\ 6(12)=k \\ 72=k \\ \text{ yeah } k \text{ is 72} \\ xy=72 \text{ replaced } k \text{ with } 72 \\ \\ \text{ now what is } x \text{ when } y=-24 \\ -24x=72 \] yep final answer is x=-3 when y=-24